OpenStudy (anonymous):

At the start of an experiment, there are 60 bacteria in a dish. The bacteria population is expected to grow at a rate of 230% each day. Which is the best prediction for the bacteria population after 5 days?

OpenStudy (anonymous):

You can use the pert formula for this equation: \[A = Pe ^{rt}\] Where P = initial amount, e = a mathematical constant, r = rate, and t = time tip: e should be on your calculator. Hope this helps!

OpenStudy (anonymous):

what would e be ?

OpenStudy (alekos):

e = 2.71828

OpenStudy (anonymous):

Thank you @alekos! I couldn't remember it off the top of my head haha

OpenStudy (mathmale):

This experiment starts with P=60 bacteria. The given growth rate is 230%, which translates into the mixed decimal fraction 2.3. The time period of interest is 5 days. You need only put this info together in the formula given you by wotseit and then evaluate it: \[A=(60 bacteria)e ^{2.3*5}\]

OpenStudy (mathmale):

What kind of calculator do you have? You'll need a calculator to evaluate the exponential function e^(2.3*5).

OpenStudy (anonymous):

uhhhh....lol i got 5922946.26 i put 60(e)^2.3 5 into my calculater ...idk what i did wrong

OpenStudy (mathmale):

I did hope you'd answer my question about the type of calculator you have. Try that calculation again. But do some preliminary work first: 2.3*5 = ?? 2.3*5 will be the exponent of your base, e.

OpenStudy (anonymous):

11.5

OpenStudy (anonymous):

and im just using a scientific calculator that is installed on my laptop

OpenStudy (anonymous):

When you are doing calculations such as this on a scientific calculator, it is important to use parenthesis: 60 x e^(2.3 *5) Otherwise, your calculator may read it as: \[60 \times e ^{2.3} \times 5\] I'm not sure if this was the problem but it helps to know anyway.

OpenStudy (mathmale):

Now take that 11.5 to be the exponent of e:\[e ^{11.5}\] Evaluate that. Then multiply the result by (60 bacteria).

OpenStudy (anonymous):

i am still getting the same answer :/

OpenStudy (mathmale):

How are you hoping I would respond?

OpenStudy (anonymous):

Can you take a screenshot of how you are putting it into your calculator? Use the prntscr button on a PC (not sure what it is for mac), paste it into paint, save it, and attach it as a file. Computer calculators are weird and I am curious how you are entering it.